OpenStudy (anonymous):

find the integral from -1 to 1 of √1+x dx

OpenStudy (raden):

integral by u-sub u = 1+x

OpenStudy (anonymous):

How do you get the 2/3?

OpenStudy (anonymous):

hey, thats wrong. dont listen to zeph

OpenStudy (anonymous):

That's wrong

OpenStudy (anonymous):

my bad

hartnn (hartnn):

did you try the u = 1+x by yourself first ?

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\sqrt{1+x} \mathrm d x\] Let u=x+1, then du=dx Find new limits of integration\[=\int\limits \sqrt{u} \mathrm d u\]

OpenStudy (anonymous):

remember square root is an exponent of 1/2

OpenStudy (isaiah.feynman):

Am I wrong? @Xeph

OpenStudy (anonymous):

i'm not sure, i got